OpenStudy (anonymous):

Speed (S) is measured in meters per second using the formula below, where d is distance traveled in meters and t is time in seconds. S =d over t If an object is moving at a speed of 9 meters per second for a time of 7 seconds, how far did the object travel?

OpenStudy (math&ing001):

S = d / t d= S * t = 9 * 7

OpenStudy (anonymous):

so the answer is 63

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

Yeah